Is it possible that a body is in accelerated motion under force acting on the body, yet no work is being done by the force? can you give an example with breif explanation.

Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

if the force is perpendicular to the instantaneous velocity, no work is done because dW=F·dr=F·vdt, best example is uniform circular motion, another example is magnetic field